# Universal relations for fundamental modes of rotating neutron stars with differential rotations

### v1

@@ Abstract
Violent astrophysical events, including core-collapse supernovae and binary neutron star mergers, can result in rotating neutron stars with diverse degrees of differential rotation. Oscillation modes of these neutron stars could be excited and emit strong gravitational waves. Detecting these modes may provide information about neutron stars, including their structures and dynamics. Hence, dynamical simulations were employed to construct relations for quantifying the oscillation mode frequency in previous studies. Specifically, linear relations for the frequencies of fundamental $l=0$ quasi-radial mode $f_{F}$ and fundamental $l=2$ quadrupolar mode $f_{{}^{2}f}$ were constructed by simulations with the Cowling approximation. Nevertheless, these relations can overestimate $f_{F}$ and underestimate $f_{{}^{2}f}$ up to $\sim 30\%$. Furthermore, it has yet to be fully studied how the degree of differential rotation affects $f_{F}$ and $f_{{}^{2}f}$. Here, for the first time, we consider both various degrees of differential rotation $\tilde{A}$ and dynamical spacetime to construct linear relations for quantifying $f_{F}$ and $f_{{}^{2}f}$. Through 2D axisymmetric simulations, we first show that both $f_{F}$ and $f_{{}^{2}f}$ scale almost linearly with the stellar compactness $M/R$ for different values of $\tilde{A}$. We also observe the quasi-linear relations for both $f_{F}$ and $f_{{}^{2}f}$ with the kinetic-to-binding energy ratio $T/|W|$ for different $\tilde{A}$ values. Finally, we constructed linear fits that can quantify $f_{F}$ and $f_{{}^{2}f}$ by $T/|W|$. Consequently, this work presented universal relations for the fundamental modes of rotating neutron stars with differential rotations in dynamical spacetime.
software: XNS (Bucciantini & Del Zanna, 2011; Pili et al., 2014, 2015, 2017; Soldateschi et al., 2020), Gmunu (Cheong et al., 2020, 2021, 2022)
1 Introduction
@@
Neutron stars with differential rotations are possible outcomes of astrophysical events, such as core-collapse supernovae and binary neutron star mergers (see e.g. Janka et al. 2007; Shibata 2016; Baiotti & Rezzolla 2017 for reviews). The degree of differential rotation of these resulting neutron stars could vary in different events. Core-collapse supernova simulations demonstrated that the ratio between central to equatorial angular velocity $\Omega_{\mathrm{c}}/\Omega_{\mathrm{e}}$ of the resulting proto-neutron stars is roughly 2 to 3 (see e.g. Yuan & Heyl 2006). On the other hand, for the hypermassive neutron stars formed in binary neutron star simulations, the angular velocity at the rotation axis could reach 1 order of magnitude higher than that at the surface (corresponds to $\Omega_{\mathrm{c}}/\Omega_{\mathrm{e}}\sim\mathcal{O}(10^{1})$) (see e.g. Baiotti et al. 2008).
@@
The neutron star remnants produced in these violent events are highly dynamical and their oscillation modes could be excited. These excited oscillation modes could be detected in gravitational wave signals (see e.g. Stergioulas et al. 2004; Kokkotas & Stergioulas 2006 for detailed discussions). Detecting these signals may provide us with information about neutron stars, such as the structure, dynamics, and equation of state (EOS) of the stars. Oscillation modes of rotating neutron stars with and without differential rotations have been extensively studied by perturbative calculations or dynamical simulations (Andersson & Kokkotas, 1998; Yoshida & Eriguchi, 1999; Kokkotas et al., 2001; Font et al., 2001; Yoshida et al., 2002; Font et al., 2002; Benhar et al., 2004; Stergioulas et al., 2004; Yoshida et al., 2005; Dimmelmeier et al., 2006; Krüger et al., 2010; Gaertig & Kokkotas, 2011; Doneva et al., 2013; Krüger & Kokkotas, 2020). In particular, Stergioulas et al. (2004) studied the axisymmetric oscillation modes of uniformly rotating neutron stars and differentially rotating neutron stars in the Cowling approximation (keeping spacetime fixed while evolving matter equations). The degree of differential rotation was fixed at $\hat{A}=A/r_{\mathrm{e}}=1$ (or equivalently $\tilde{A}=(A/r_{\mathrm{e}})^{-1}=1$ in this work) to mimick the rotational profiles of proto-neutron stars, where $r_{\mathrm{e}}$ is the equatorial radius of the star. They found that the frequencies of both fundamental $l=0$ quasi-radial mode $f_{F}$ and fundamental $l=2$ quadrupolar mode $f_{{}^{2}f}$ of rotating neutron stars scale almost linearly with the kinetic-to-binding energy ratio $T/|W|$. Then, by assuming a fixed relative differences between the actual frequencies and those obtained in their simulations in the Cowling approximation, Stergioulas et al. (2004) also constructed linear relations as functions of $T/|W|$ for predicting the actual frequencies of $f_{F}$ and $f_{{}^{2}f}$ of rotating neutron stars. Dimmelmeier et al. (2006) later re-examined this problem by including the effect of dynamical spacetime. They demonstrated that simulations using the Cowling approximation can overestimate oscillation mode frequency up to a factor of 2. They also found that the relative differences between the actual frequencies and those obtained in the Cowling approximation significantly change as $T/|W|$ increases. As a consequence, applying linear relations derived from Stergioulas et al. (2004) to predict actual frequencies can lead to an overestimate of $f_{F}$ and an underestimate of $f_{{}^{2}f}$ up to $\sim 30\%$. Yet, Dimmelmeier et al. (2006) also considered differentially rotating neutron stars with $\hat{A}=1$ only. For a better understanding of neutron star oscillations under various astrophysical scenarios, a study that takes into account various degrees of differential rotation and dynamical spacetime is required.
@@
An ideal tool for conducting such a study would be dynamical simulations capable of examining a large number of stellar models in a reasonable amount of time and with minimal computational resources. This kind of simulation can be performed using Gmunu (Cheong et al., 2020, 2021, 2022), a new general-relativistic magnetohydrodynamics code. In Gmunu, simulations can be performed in multiple dimensions (1D, 2D, and 3D) as well as coordinate systems (cartesian, cylindrical, and spherical) using a block-based adaptive mesh refinement (AMR) module. These properties enable imposing symmetries to reduce the dimensionality and computational costs of different problems considerably. Besides, using the multigrid method, Gmunu also solves the elliptic metric equations in the conformally flat condition (CFC) approximation efficiently and robustly. All the features mentioned above make Gmunu an appropriate tool for studying problems related to a single neutron star (e.g. Ng et al. 2021; Leung et al. 2022; Yip et al. 2023a, b, 2023).
@@
In this work, for the first time, we take into account both various degrees of differential rotation and dynamical spacetime to construct linear relations quantifying $f_{F}$ and $f_{{}^{2}f}$. Specifically, we first construct initial neutron star models by the open-sourced code XNS (Bucciantini & Del Zanna, 2011; Pili et al., 2014, 2015, 2017; Soldateschi et al., 2020). The equilibrium models are then perturbed and evolved in dynamical spacetime using Gmunu. The details of the initial neutron star models and evolutions are described in Section 2. After that, we examine how the frequencies of fundamental modes vary with the compactness $M/R$ and the kinetic-to-binding energy ratio $T/|W|$ of neutron stars in Sections 3 and 4 respectively. Finally, we construct linear relations as functions of $T/|W|$ for quantifying $f_{F}$ and $f_{{}^{2}f}$ of rotating neutron stars with differential rotations and provide the conclusions in Section 5.
@@
2 Numerical methods
2.1 Initial neutron star models
We compute the neutron star equilibrium models in axisymmetry by the open-sourced code XNS (Bucciantini & Del Zanna, 2011; Pili et al., 2014, 2015, 2017; Soldateschi et al., 2020). Our simulations are conducted using these models as initial data.
@@ 2.1 Initial neutron star models
The $j$-constant law introduced by Komatsu et al. (1989a, b) has been widely used to model differentially rotating neutron stars (see e.g. Baumgarte et al. 2000; Morrison et al. 2004; Kaplan et al. 2014 )
@@
where $j$ is the relativistic specific angular momentum, $A$ is a parameter to control the degree of differential rotation, $\Omega$ is the angular velocity and $\Omega_{\mathrm{c}}$ is the central angular velocity. Following previous studies (e.g. Krüger et al. 2010), we adopt the parameter $\tilde{A}=(A/r_{\mathrm{e}})^{-1}$ to quantify the degree of differential rotation, where $r_{\mathrm{e}}$ is the equatorial radius of the equilibrium model. Greater $\tilde{A}$ corresponds to a higher degree of differential rotation and the star undergoes uniform rotation when $\tilde{A}=0$. We constructed equilibrium models in 5 sequences with $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$ to investigate the effect of the degree of differential rotation. As we do not intend to investigate neutron stars with different masses in this work, we adopt a fixed baryonic mass $M_{0}=1.506$ for all models. 3 neutron stars with high masses have been observed recently: J0348+0432 at $M=2.01\pm 0.04$ (Antoniadis et al., 2013), PSR J0740+6620 at $M=2.08\pm 0.07$ (Fonseca et al., 2021), and PSR J0952-0607 at $M=2.35\pm 0.17$ (Romani et al., 2022). In light of these observations, a neutron star should have a maximum mass of at least $M=2.0$. Accordingly, the masses of our models are within the observational constraints on the maximum mass of a neutron star. Besides, the adopted baryonic mass $M_{0}=1.506$ is identical to that for the models in the previous study by Dimmelmeier et al. (2006), enabling comparison with that study. The detailed properties of the equilibrium models are summarized in Appendix A.
2.2 Evolutions
@@
Using the new general relativistic magnetohydrodynamics code Gmunu (Cheong et al., 2020, 2021, 2022), we evolve the stellar models in dynamical spacetime. The models evolved over a time of 20 ms with the polytropic equation of state $P=K\rho^{\gamma}$, under the same conditions as equilibrium models (i.e. $K=100$ and $\gamma=2$).
@@
2D ideal general-relativistic hydrodynamics simulations are conducted in axisymmetry with respect to the $z$-axis and equatorial symmetry using spherical coordinates $(r,\theta)$. The computational domain covers $0\leq r\leq 60$, $20\leq\theta\leq\pi/$ 2, with the base grid resolution $N_{r}\times N_{\theta}=64\times 16$ and allowing 4 AMR levels (effective resolution = $512\times 128$). The refinement criteria of AMR is equivalent to that in Cheong et al. (2021); Leung et al. (2022); Yip et al. (2023a, b). TVDLF approximate Riemann solver (Tóth & Odstrčil, 1996) , 3rd-order reconstruction method PPM (Colella & Woodward, 1984) and 3rd-order accurate SSPRK3 time integrator (Shu & Osher, 1988) are employed in our simulations. Outside the star, there is an artificial atmosphere with rest-mass density $\rho_{\mathrm{atm}}\sim 10^{-10}\rho_{\mathrm{c}}$.
2.3 Initial perturbations
@@
Following Dimmelmeier et al. (2006), we excite the fundamental modes by applying the following initial fluid perturbations on the equilibrium models.
@@
First, we use the $l=0$ perturbation on the $r$-component of the three-velocity field $v^{r}$ to excite the fundamental $l=0$ quasi-radial mode (i.e. ${F}$-mode),
@@
Fundamental modes are extracted by performing Fourier transforms of $v^{r}$ and $v^{\theta}$ at $r=3$, $\theta=\pi/4$, ensuring the extraction position is within the star (see Dimmelmeier et al. 2006 for more details).
@@
3 Fundamental mode frequencies against stellar compactness
@@
As shown by Hartle & Friedman (1975), the frequencies of the fundamental modes are related to the stellar compactness of neutron stars. To determine whether the degree of differential rotation affects this relation, we plot the fundamental mode frequency $f$ against the stellar compactness $M/R$ in Fig. 1, where $M$ is the stellar mass and $R$ is the circumferential radius. The data points are arranged into 5 sequences with $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$, where $\tilde{A}$ is the degree of differential rotation. $\tilde{A}=0.0$ refers to uniformly rotating cases. Data points connected by solid lines represent the data for the frequency of fundamental $l=0$ quasi-radial mode $f_{F}$ while the points connected by dashed lines denote the data for the frequency of fundamental $l=2$ quadrupolar mode $f_{{}^{2}f}$. The frequency $f$ of both fundamental modes increases with $M/R$ almost linearly. Slope of $f(M/R)$ for both fundamental modes changes only slightly with $\tilde{A}$, which is within our simulation error. Hence, we observe a quasi-linear relation between fundamental mode frequency and stellar compactness for rotating neutron stars with different degrees of differential rotation.
@@
Figure 1: Plot of fundamental mode frequency $f$ against stellar compactness $M/R$, where $M$ is the stellar mass and $R$ is the circumferential radius. The data points are arranged into 5 sequences with $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$, where $\tilde{A}$ is the degree of differential rotation. $\tilde{A}=0.0$ refers to uniformly rotating models. Data points connected by solid lines denote the data for the frequency of fundamental $l=0$ quasi-radial mode $f_{F}$ while the points connected by dashed lines denote the data for the frequency of fundamental $l=2$ quadrupolar mode $f_{{}^{2}f}$. Both $f_{F}$ and $f_{{}^{2}f}$ increase with $M/R$ in an approximately linear way. The slopes varies only slightly with $\tilde{A}$, which is within our simulation error. Hence, this demonstrates a quasi-linear relation between fundamental mode frequency and stellar compactness for rotating neutron stars with different degrees of differential rotation.
@@
4 Fundamental mode frequencies against kinetic-to-binding energy ratio

### v2

# Fundamental modes of rotating neutron stars with various degrees of differential rotation in dynamical spacetimes

@@ Abstract
Violent astrophysical events, including core-collapse supernovae and binary neutron star mergers, can result in rotating neutron stars with diverse degrees of differential rotation. Oscillation modes of these neutron stars could be excited and emit strong gravitational waves. Detecting these modes may provide information about neutron stars, including their structures and dynamics. Hence, dynamical simulations were employed to construct relations for quantifying the oscillation mode frequency in previous studies. Specifically, linear relations for the frequencies of fundamental $l=0$ quasi-radial mode $f_{F}$ and fundamental $l=2$ quadrupolar mode $f_{{}^{2}f}$ were constructed by simulations with the Cowling approximation. Nevertheless, these relations can overestimate $f_{F}$ and underestimate $f_{{}^{2}f}$ up to $\sim 30\%$. Furthermore, it has yet to be fully studied how the degree of differential rotation affects $f_{F}$ and $f_{{}^{2}f}$. Here, for the first time, we consider both various degrees of differential rotation $\tilde{A}$ and dynamical spacetime to construct linear relations for quantifying $f_{F}$ and $f_{{}^{2}f}$. Through 2D axisymmetric simulations, we first show that both $f_{F}$ and $f_{{}^{2}f}$ scale almost linearly with the stellar compactness $M/R$ for different values of $\tilde{A}$. We also observe the quasi-linear relations for both $f_{F}$ and $f_{{}^{2}f}$ with the kinetic-to-binding energy ratio $T/|W|$ for different $\tilde{A}$ values. Finally, we constructed linear fits that can quantify $f_{F}$ and $f_{{}^{2}f}$ by $T/|W|$. Consequently, this work updated the relations for the fundamental modes of rotating neutron stars with differential rotations in dynamical spacetime.
preprint: APS/123-QED
I Introduction
@@
Neutron stars with differential rotations are possible outcomes of astrophysical events, such as core-collapse supernovae and binary neutron star mergers (see e.g. [1, 2, 3] for reviews). The degree of differential rotation of these resulting neutron stars could vary in different events. Core-collapse supernova simulations demonstrated that the ratio between central to equatorial angular velocity $\Omega_{\mathrm{c}}/\Omega_{\mathrm{e}}$ of the resulting proto-neutron stars is roughly 2 to 3 (see e.g. [4]). On the other hand, for the hypermassive neutron stars formed in binary neutron star simulations, the angular velocity at the rotation axis could reach 1 order of magnitude higher than that at the surface (corresponds to $\Omega_{\mathrm{c}}/\Omega_{\mathrm{e}}\sim\mathcal{O}(10^{1})$) (see e.g. [5]).
@@
The neutron star remnants produced in these violent events are highly dynamical and their oscillation modes could be excited. These excited oscillation modes could be detected in gravitational wave signals (see e.g. [6, 7] for detailed discussions). Detecting these signals may provide us with information about neutron stars, such as the structure, dynamics, and equation of state (EOS) of the stars (see e.g. [8] for a review of oscillations modes). Oscillation modes of rotating neutron stars with and without differential rotations have been extensively studied by perturbative calculations or dynamical simulations [9, 10, 11, 12, 13, 14, 15, 6, 16, 17, 18, 19, 20, 21]. In particular, [6] studied the axisymmetric oscillation modes of uniformly rotating neutron stars and differentially rotating neutron stars in the Cowling approximation (keeping spacetime fixed while evolving matter equations). The degree of differential rotation was fixed at $\hat{A}=A/r_{\mathrm{e}}=1$ (or equivalently $\tilde{A}=(A/r_{\mathrm{e}})^{-1}=1$ in this work) to mimick the rotational profiles of proto-neutron stars, where $r_{\mathrm{e}}$ is the equatorial radius of the star. They found that the frequencies of both fundamental $l=0$ quasi-radial mode $f_{F}$ and fundamental $l=2$ quadrupolar mode $f_{{}^{2}f}$ of rotating neutron stars scale almost linearly with the kinetic-to-binding energy ratio $T/|W|$. Then, by assuming a fixed relative differences between the actual frequencies and those obtained in their simulations in the Cowling approximation, [6] also constructed linear relations as functions of $T/|W|$ for predicting the actual frequencies of $f_{F}$ and $f_{{}^{2}f}$ of rotating neutron stars. [17] later re-examined this problem by including the effect of dynamical spacetime. They demonstrated that simulations using the Cowling approximation can overestimate oscillation mode frequency up to a factor of 2. They also found that the relative differences between the actual frequencies and those obtained in the Cowling approximation significantly change as $T/|W|$ increases. As a consequence, applying linear relations derived from [6] to predict actual frequencies can lead to an overestimate of $f_{F}$ and an underestimate of $f_{{}^{2}f}$ up to $\sim 30\%$. Yet, [17] also considered differentially rotating neutron stars with $\hat{A}=1$ only. For a better understanding of neutron star oscillations under various astrophysical scenarios, a study that takes into account various degrees of differential rotation and dynamical spacetime is required.
@@
An ideal tool for conducting such a study would be dynamical simulations capable of examining a large number of stellar models in a reasonable amount of time and with minimal computational resources. This kind of simulation can be performed using Gmunu [22, 23, 24], a new general-relativistic magnetohydrodynamics code. In Gmunu, simulations can be performed in multiple dimensions (1D, 2D, and 3D) as well as coordinate systems (cartesian, cylindrical, and spherical) using a block-based adaptive mesh refinement (AMR) module. This flexibility allows users to select the most appropriate dimensionality and coordinate system for each problem, while also imposing symmetries to further reduce computational costs where possible. For studies focused on axisymmetric ($m=0$) oscillation modes (e.g. $F$-mode and $^{2}f-$mode investigated by [6, 17]), 2D axisymmetric simulations performed in this study are particularly well-suited. However, it is important to note that non-axisymmetric modes, such as the $m=1$ and $m=2$ modes, can arise and become significant in binary neutron star remnants and proto-neutron stars formed during core-collapse supernovae (see e.g. [25]). As a result, the assumption of 2D axisymmetry in this study hinders the capture of these non-axisymmetric modes in the simulations. Investigating these modes would require 3D simulations without imposing axisymmetry. Besides, using the multigrid method, Gmunu also solves the elliptic metric equations in the conformally flat condition (CFC) approximation efficiently and robustly. All the features mentioned above make Gmunu an appropriate tool for studying problems related to a single neutron star (e.g. [26, 27, 28, 29, 30]).
@@
In this work, for the first time, we take into account both various degrees of differential rotation and dynamical spacetime to construct linear relations quantifying $f_{F}$ and $f_{{}^{2}f}$. Specifically, we first construct initial neutron star models by the open-sourced code XNS [31, 32, 33, 34, 35]. A variety of differential rotation laws have been proposed in the literature to describe the differential rotation of protoneutron stars and binary neutron star merger remnants (see e.g [36, 37, 38, 39]). Similarly, realistic equations of state, including finite temperature effects, have been used to model these neutron star remnants (see e.g [3, 40, 41, 42]). However, the primary focus of this study is on the effects of degrees of differential rotation and dynamical spacetime rather than the effects coming from the details of the variety of rotation laws and equations of state discussed in the literature. Therefore, following previous works [6, 17], we adopt a simplified case of the $j$-constant differential rotation law [43, 44] and a polytropic equation of state, without considering finite-temperature effects or realistic equations of state, to construct the initial neutron star models. While this approach allows us to isolate the qualitative impact of differential rotation and dynamical spacetime, these simplifying assumptions can lead to quantitative deviations in the computed mode frequencies. Thus, more realistic equations of state and rotation laws are necessary for a more accurate determination of the mode frequencies in scenarios such as protoneutron stars and binary neutron star merger remnants. These equilibrium models are then perturbed and evolved in dynamical spacetime using Gmunu. The details of the initial neutron star models and evolutions are described in Section II. After that, we examine how the frequencies of fundamental modes vary with the compactness $M/R$ and the kinetic-to-binding energy ratio $T/|W|$ of neutron stars in Sections III and IV respectively. Finally, we construct linear relations as functions of $T/|W|$ for quantifying $f_{F}$ and $f_{{}^{2}f}$ of rotating neutron stars with differential rotations and provide the conclusions in Section V.
@@
II Numerical methods
II.1 Initial neutron star models
We compute the neutron star equilibrium models in axisymmetry by the open-sourced code XNS [31, 32, 33, 34, 35]. Our simulations are conducted using these models as initial data.
@@ II.1 Initial neutron star models
The $j$-constant law introduced by [43, 44] has been widely used to model differentially rotating neutron stars (see e.g. [45, 46, 47] )
@@
where $j$ is the relativistic specific angular momentum, $A$ is a parameter to control the degree of differential rotation, $\Omega$ is the angular velocity and $\Omega_{\mathrm{c}}$ is the central angular velocity. Following previous studies (e.g. [18]), we adopt the parameter $\tilde{A}=(A/r_{\mathrm{e}})^{-1}$ to quantify the degree of differential rotation, where $r_{\mathrm{e}}$ is the equatorial radius of the equilibrium model. Greater $\tilde{A}$ corresponds to a higher degree of differential rotation and the star undergoes uniform rotation when $\tilde{A}=0$. We constructed equilibrium models in 5 sequences with $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$ to investigate the effect of the degree of differential rotation. As we do not intend to investigate neutron stars with different masses in this work, we adopt a fixed baryonic mass $M_{0}=1.506$ for all models. 3 neutron stars with high masses have been observed recently: J0348+0432 at $M=2.01\pm 0.04$ [48], PSR J0740+6620 at $M=2.08\pm 0.07$ [49], and PSR J0952-0607 at $M=2.35\pm 0.17$ [50]. In light of these observations, a neutron star should have a maximum mass of at least $M=2.0$. Accordingly, the masses of our models are within the observational constraints on the maximum mass of a neutron star. Besides, the adopted baryonic mass $M_{0}=1.506$ is identical to that for the models in the previous study by [17], enabling comparison with that study. The detailed properties of the equilibrium models are summarized in Appendix A.
II.2 Evolutions
@@
Using the new general relativistic magnetohydrodynamics code Gmunu [22, 23, 24], we evolve the stellar models in dynamical spacetime. The models evolved over a time of 20 ms with the polytropic equation of state $P=K\rho^{\gamma}$, under the same conditions as equilibrium models (i.e. $K=100$ and $\gamma=2$).
@@
2D ideal general-relativistic hydrodynamics simulations are conducted in axisymmetry with respect to the $z$-axis and equatorial symmetry using spherical coordinates $(r,\theta)$. The computational domain covers $0\leq r\leq 60$, $0\leq\theta\leq\pi/2$, with the base grid resolution $N_{r}\times N_{\theta}=64\times 16$ and allowing 4 AMR levels (effective resolution = $512\times 128$). The refinement criteria of AMR is equivalent to that in [23, 27, 28, 29]. TVDLF approximate Riemann solver [51] , 3rd-order reconstruction method PPM [52] and 3rd-order accurate SSPRK3 time integrator [53] are employed in our simulations. Outside the star, there is an artificial atmosphere with rest-mass density $\rho_{\mathrm{atm}}\sim 10^{-10}\rho_{\mathrm{c}}$.
II.3 Initial perturbations
@@
Following [17], we excite the fundamental modes by applying the following initial fluid perturbations on the equilibrium models.
@@
First, we use the $l=0$ perturbation on the $r$-component of the three-velocity field $v^{r}$ to excite the fundamental $l=0$ quasi-radial mode (i.e. F{F}-mode),
@@
Fundamental modes are extracted by performing Fourier transforms of $v^{r}$ and $v^{\theta}$ at $r=3$, $\theta=\pi/4$, ensuring the extraction position is within the star (see Appendix B or [17] for more details).
@@
III Fundamental mode frequencies against stellar compactness and radius ratio
@@
As the rotation rate of a neutron star increases, both the stellar compactness and radius ratio decrease, given that the baryonic mass remains constant (see e.g. [17]). [54] have also shown that the frequencies of the fundamental modes are related to the stellar compactness of neutron stars. To determine how how the fundamental mode frequency is related to both the stellar compactness and radius ratio in rotating neutron stars with different degrees of differential rotation, we plot the fundamental mode frequency $f$ against the stellar compactness $M/R$ (left panel) and radius ratio $r_{\mathrm{p}}/r_{\mathrm{e}}$ (right panel) in Fig. 1, where $M$ is the stellar mass, $R$ is the circumferential radius, $r_{\mathrm{p}}$ is the polar radius and $r_{\mathrm{e}}$ is the equatorial radius. The data points are arranged into 5 sequences with $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$, where $\tilde{A}$ is the degree of differential rotation. $\tilde{A}=0.0$ refers to uniformly rotating cases. Data points connected by solid lines represent the data for the frequency of fundamental $l=0$ quasi-radial mode $f_{F}$ while the points connected by dashed lines denote the data for the frequency of fundamental $l=2$ quadrupolar mode $f_{{}^{2}f}$. The frequency $f$ of both fundamental modes increases with $M/R$ and $r_{\mathrm{p}}/r_{\mathrm{e}}$ almost linearly. Slopes of $f(M/R)$ and $f(r_{\mathrm{p}}/r_{\mathrm{e}})$ for both fundamental modes changes only slightly with $\tilde{A}$. Hence, we observe a quasi-linear relation between fundamental mode frequency, stellar compactness and radius ratio for rotating neutron stars with different degrees of differential rotation.
@@
Figure 1: Plot of fundamental mode frequency $f$ against stellar compactness $M/R$ (left panel) and radius ratio $r_{\mathrm{p}}/r_{\mathrm{e}}$ (right panel), where $M$ is the stellar mass, $R$ is the circumferential radius, $r_{\mathrm{p}}$ is the polar radius and $r_{\mathrm{e}}$ is the equatorial radius. The data points are arranged into 5 sequences with $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$, where $\tilde{A}$ is the degree of differential rotation. $\tilde{A}=0.0$ refers to uniformly rotating models. Data points connected by solid lines denote the data for the frequency of fundamental $l=0$ quasi-radial mode $f_{F}$ while the points connected by dashed lines denote the data for the frequency of fundamental $l=2$ quadrupolar mode $f_{{}^{2}f}$. Both $f_{F}$ and $f_{{}^{2}f}$ increase approximately linearly with $M/R$ and $r_{\mathrm{p}}/r_{\mathrm{e}}$, with the slopes varying only slightly as $\tilde{A}$ changes. Hence, this demonstrates a quasi-linear relation between fundamental mode frequency, stellar compactness and radius ratio for rotating neutron stars with different degrees of differential rotation.
@@
IV Fundamental mode frequencies against kinetic-to-binding energy ratio